OpenStudy (anonymous):

how is the the ''12 less than x is four'' written as equation?

OpenStudy (anonymous):

Four times the lesser of two consecutive even integers is 12 less than twice the greater number. Find the integers. .. let the numbers be x & x+2 4x=2(x+2)-12 4x=2x+4-12 4x-2x=-8 2x=-8 /2 x=-4 the numbers are -4 & -2

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

anytime